OpenStudy (anonymous):

Solve using the elimination method. -3x + 4y = 1 -6x + 6y = -3 i got as far as eliminating x so I have 10y=-3 ?

OpenStudy (anonymous):

I mean 14y=-3 is this correct?

OpenStudy (anonymous):

(-3x+4y=1) <<<Multiply this by -2 to elimiinate x -6x+6y=-3

OpenStudy (anonymous):

right so therefore the problem becomes -2y=1

OpenStudy (anonymous):

No!, You're left with -2y=-5

OpenStudy (anonymous):

oh ok ad both sides gotcha

OpenStudy (anonymous):

then it would be y=-5/2 is this correct?

OpenStudy (anonymous):

y=5/2, dividing a negative by a negative yields a positive

OpenStudy (anonymous):

ok then I just plug 5/2 into both equations to solve for x correct?

OpenStudy (anonymous):

Plug it into 1 will work

OpenStudy (anonymous):

sorry I am really trying to figure out how to do this not just the answer. I have been working on it for a couple of days now and really am not understanding it

OpenStudy (anonymous):

No big deal, you'll get it

OpenStudy (anonymous):

Thank you for all of your help

OpenStudy (anonymous):

NP